OpenStudy (anonymous):

Simplify. Express the product as a rational exponent. x^1/3 times x^1/3

OpenStudy (anonymous):

a^b * a^c = a^(b + c) a^b / a^c = a^(b - c) (x^(1/3)) * (x^(1/3)) => x^(1/3 + 1/3) => x^(2/3)

OpenStudy (anonymous):

you're a life saver @aballer101

OpenStudy (anonymous):

i.e add \[\frac{1}{3}+\frac{1}{3}\]

OpenStudy (anonymous):

well thats simple to remember haaha @satellite73

OpenStudy (anonymous):

not too bad, no

OpenStudy (anonymous):

ur welcome! :)